6=-4+h
-7+d=-3
9=m+11
f+(-9)=-19

Answer by elima(1433)

6=-4+h
Add 4 to both sides
6+4=-4+h=4
10=h
-7+d=-3
add 7 to both sides;
-7+d+7=3+7
d=10
9=m+11
Subtract both sides by 11;
9-11=m+11-11
-2=m
f+(-9)=-19
add 9 to both sides;
f+(-9)+9=-19+9
f=-10
=)